OpenStudy (anonymous):

Simplify \(\frac{cos^{2}a}{1-tana} + \frac{sin^{2}a}{1-cota}\)

OpenStudy (loser66):

I saw you post it twice, I will guide you how to do, but you WORK, not me, OK?

OpenStudy (anonymous):

yeah

OpenStudy (loser66):

break it down, the first term: I let a aside because all of them respect to a, right? \[\dfrac{cos^2}{1-tan}=\dfrac{cos^2}{1-\dfrac{sin}{cos}}\] .Simplify it, what do you get?

OpenStudy (anonymous):

thank you :) I've solved it

OpenStudy (loser66):

ok, good luck